OpenStudy (credmond):

Use the product rule to find the derivative of y= (2x-3)(sqrt x - 1)

OpenStudy (nincompoop):

product rule states y' = u'v + uv'

OpenStudy (nincompoop):

but you're probably going to need more than the product rule in the problem you're solving it also requires you to use chain rule because of the sqrt(x-1)

OpenStudy (nincompoop):

I've given you the value of u' already now, it's your turn to finish the problem

OpenStudy (credmond):

Thanks for the help! I really appreciate it! :) But I'm pretty lost.

OpenStudy (anonymous):

open the brackets first... den apply chain rule.. its easy...

OpenStudy (anonymous):

product rule^

OpenStudy (credmond):

Okay, actually I followed that first step you did. I just don't know where to go next

OpenStudy (anonymous):

study product rule and chain rule.. i can tell the answer but it wont b useful for u... practice some product rule and chain rule ques.

OpenStudy (credmond):

The problem is that we haven't learned about the chain rule yet! :/ That's the next section we are doing though! So I'm having a hard time proceeding with this question.

OpenStudy (nincompoop):

chain rule is a combination of rules really

OpenStudy (nincompoop):

you follow the sequence like with higher order of operations example sqrt(2x^3 + 1) can be rewritten into (2x^3 +1)^(1/2) first apply exponents outside - the ^(1/2), this means you need a power rule 1/2 (2x^3+1) then you obtain the derivative of whatever is inside the radical - the 2x^3 + 1 6x^2 now combine them by means of multiplication 1/2 (2x^3+1) * 6x^2 simplify further and you're done

OpenStudy (nincompoop):

so go back to what the product rule says y' = u'v + uv' y= (2x-3)(sqrt x - 1), y'=? substitute the following (2x-3) = u, then u' = 2 sqrt(x-1) = v, then v' = 1/2sqrt(x-1)^2

OpenStudy (nincompoop):

I am not sure if you're using a Leibniz notation, and sorry for not using it :)

OpenStudy (credmond):

Ohhhhh. Okay gotcha! Sometimes I just have to have people walk me through the problem every now and then. I really appreciate your help! :) I'll take a look at that link as well!